# I2S-TFCKD: Intra-Inter Set Knowledge Distillation with Time-Frequency Calibration for Speech Enhancement

### v1

@@ Abstract
In recent years, complexity compression of neural network (NN)-based speech enhancement (SE) models has gradually attracted the attention of researchers, especially in scenarios with limited hardware resources or strict latency requirements. The main difficulties and challenges lie in achieving a balance between complexity and performance according to the characteristics of the task. In this paper, we propose an intra-inter set knowledge distillation (KD) framework with time-frequency calibration (I2S-TFCKD) for SE. Different from previous distillation strategies for SE, the proposed framework fully utilizes the time-frequency differential information of speech while promoting global knowledge flow. Firstly, we propose a multi-layer interactive distillation based on dual-stream time-frequency cross-calibration, which calculates the teacher-student similarity calibration weights in the time and frequency domains respectively and performs cross-weighting, thus enabling refined allocation of distillation contributions across different layers according to speech characteristics. Secondly, we construct a collaborative distillation paradigm for intra-set and inter-set correlations. Within a correlated set, multi-layer teacher-student features are pairwise matched for calibrated distillation. Subsequently, we generate representative features from each correlated set through residual fusion to form the fused feature set that enables inter-set knowledge interaction. The proposed distillation strategy is applied to the dual-path dilated convolutional recurrent network (DPDCRN) that ranked first in the SE track of the L3DAS23 challenge. To evaluate the effectiveness of I2S-TFCKD, we conduct experiments on both single-channel and multi-channel SE datasets. Objective evaluations demonstrate that the proposed KD strategy consistently and effectively improves the performance of the low-complexity student model and outperforms other distillation schemes.
Index Terms:
speech enhancement, knowledge distillation, time-frequency calibration, intra-inter set correlations, residual fusion.
publicationid: pubid: $0$–$00$ 2025 IEEE
I Introduction
@@
Speech enhancement (SE) aims to remove background noises and recover clean speech from noisy mixtures. As a perception front-end algorithm, its applications have already penetrated into diverse domains including such as intelligent terminals, industrial manufacturing, medical health, etc. [1]. Traditional SE methods based on statistical signal processing [2] estimate the noise spectrum from noisy speech for suppression, requiring low computational resources. However, these algorithms typically rely on stationarity assumptions, which fundamentally constrain their performance in handling non-stationary noises encountered in real-world scenarios [3]. With the recent emergence of data-driven deep learning approaches, SE models have begun leveraging massive training data to characterize both speech signals, noise components, and their nonlinear relationships. Capitalizing on the representational learning capabilities of deep neural networks (DNNs), deep learning-based SE methods have achieved superior noise suppression performances and have attained state-of-the-art results in international SE challenges [4], gradually becoming the mainstream SE solutions.
@@
Model compression is an effective solution, including pruning [7], quantization [8], and knowledge distillation (KD) [9]. Stamenovic et al. [10] specifically investigate pruning methodologies tailored for SE frameworks. Tan et al. [11] reduce the complexity of SE networks using three different techniques: sparse regularization, iterative pruning, and clustering-based quantization. However, pruning and quantization still heavily depend on the original model structure and hardware support. In contrast, KD transfers knowledge from a powerful teacher to a compact student, offering better generality and flexibility—thus becoming the technical path chosen in this paper. Actually, many researchers have already focused on distillation methods for SE models. Kim et al. [12] uses the teacher’s pseudo target to train the student model, enabling a zero-shot learning procedure. Thakker et al. [13] integrate KD with multi-task learning losses from speech recognition to achieve SE model compression. These studies primarily focus on minimizing output-level discrepancies between teachers and students while neglecting intermediate feature representations. Recently, Wan et al. [14] propose a hierarchical attention-based distillation framework for SE, exploring the utilization of intermediate representations. Similarly, Nathoo et al. [15] introduce a fine-grained similarity-preserving KD loss, which aims to match the student’s intra-activation Gram matrices to that of the teacher. Knowledge transfer of intermediate model representations has gradually attracted researchers’ attention, with the key challenge lying in the SE-specific distillation design.
@@
In our previous work [16], we proposed a unified residual fusion probabilistic KD framework for SE, which leverages cross-layer supervision to guide information transfer between teacher and student while introducing a frame weighting probabilistic distillation loss to perceive key frame information. In this paper, we design an intra-inter set KD with time-frequency calibration (I2S-TFCKD) framework to further optimize the cross-layer supervision distillation paradigm. On the one hand, we introduce temporal and spectral domain cross-computation to derive multi-layer distillation calibration weights, leveraging time-frequency characteristics for targeted distillation optimization. On the other hand, unlike previous approaches that conduct cross-layer interactions within isolated correlation sets, we propose collaborative distillation across intra-set and inter-set correlations, effectively promoting knowledge flow throughout the model architecture. The backbone network for distillation employs our previously proposed dual-path dilated convolutional recurrent network (DPDCRN) model [17], which won the SE track of the L3DAS23 challenge. The main contributions of this work are as follows:
  • We propose multi-layer time-frequency cross-calibration distillation, which independently computes teacher-student similarity calibration weights along temporal and spectral dimensions for cross-weighting, enabling refined allocation of distillation contributions across different layers.
@@
  • We design a collaborative distillation paradigm for intra-set and inter-set correlations. First, we partition the SE model structure into multiple correlation sets for intra-set multi-layer teacher-student matching. Subsequently, representative features from each correlated set are generated through residual recursive fusion, forming the fused feature set that enables inter-set knowledge interaction.
@@
  • Extensive evaluations on public SE benchmarks demonstrate that our proposed I2S-TFCKD framework for SE surpasses existing distillation methods and enabling the student model to achieve competitive enhancement results with low computational complexity.
II Related Work
In this section, we provide an overview of related works on distillation methodologies for SE, particularly focusing on feature distillation-related research.
II-A Knowledge Distillation Strategy
KD originated from the pioneering work of Hinton et al. [18], where the student model learns from both ground-truth labels and soft labels provided by the teacher model. Building upon this, subsequent research focuses on exploring effective regularization methods to balance soft and hard labels. In recent years, feature distillation has gradually come into the spotlight, with numerous studies demonstrating that distilling intermediate feature representations exhibits stronger generalization ability and more promising results compared to soft-label distillation. Niu et al. [19] transform the feature maps of different levels into a frequency-enhanced feature space, which highlights the low-frequency features for both the teacher and student networks. Cross-image relational knowledge extraction is proposed in [20], leveraging pixel-to-pixel and pixel-to-region distillation to fully utilize the structural relationships between images. These methods primarily utilize single-layer features for knowledge transfer, while recent advances explore multi-layer and multi-stage feature distillation frameworks. A cross-layer knowledge review mechanism is investigated in [21], which integrates multi-layer teacher features to guide the student. Zhao et al. [22] developed a dense KD in all aspects manner, using the dynamic class-specific pseudo-labeling strategy to distill intermediate features and output logits collaboratively with more emphasis on semantic-invariant knowledge inheritance. Overall, feature distillation can transfer fine-grained knowledge such as edge, texture, and semantic features, garnering increasing research attention. Cross-layer feature distillation transcends traditional layer-wise correspondence constraints, allowing a student to simultaneously learn low-level detail features and high-level abstract representations from teachers. However, this paradigm also introduces challenges in representation alignment, making it a worthwhile direction for in-depth research.
II-B Distillation for Speech Enhancement
@@
Most existing distillation works discussed above have been primarily applied to the computer vision and natural language processing domains. In the acoustic field, KD methods are initially adopted for classification tasks such as acoustic scene classification [23] and automatic speech recognition [24]. Subsequently, KD for regression tasks like SE has gradually emerged. Early distillation studies for SE models primarily focus on minimizing output-level discrepancies between teachers and students, exemplified by elite sub-band distillation [25]. These works aim to align student models with teacher patterns at the learning objective level but overlook the utilization of intermediate representations containing richer information. Feature distillation holds greater potential for SE models as their architectures and functionalities grow increasingly complex. A common challenge in representation-based distillation arises from dimension mismatch between teacher and student features. To address this, our previous work [26] first proposes a cross-layer similarity distillation framework for SE models, which aligns teacher-student features via similarity loss. Similarly, multiple studies have adopted multi-layer distillation strategies to compress SE models, including multi-view attention transfer distillation [27] and two-step fine-grained similarity-preserving KD [15]. It is worth noting that recent studies have begun to focus on leveraging the inherent characteristics of the SE task to improve distillation frameworks. Dynamic Frequency Adaptive Distillation [28] dynamically evaluates model outputs and adjusts distillation targets according to the requirements of different speech frequency bands. An independent attention transfer mechanism is designed for temporal and channel dimensions in [29] to facilitate rich knowledge transfer across different structured models. However, existing studies have not fully exploited the time-frequency characteristics of speech in designing feature distillation frameworks. Additionally, cross-layer distillation frameworks such as [15, 16] do not consider the global propagation of knowledge. To address these limitations, this work mainly includes two innovative improvements: the time-frequency calibration strategy for distillation and the intra-inter set global knowledge transfer mechanism.
III Methodology
III-A Teacher and Student Network Architecture
@@
Figure 2: Overall architecture of the I2S-TFCKD framework. We present the detailed process of intra-inter set distillation. Fig. 2(a) shows the time-frequency cross-calibration knowledge transfer and the recursive feature fusion across different layers within a single correlated set. Fig. 2(b) demonstrates inter-set distillation achieved via the fused feature set among various correlated sets. Fig. 2(c) visualizes the process of generating fused features by utilizing the features of the current layer and the inherited recursive features.
@@
In our previous work [16], we designed a residual fusion distillation framework for SE based on the DPDCRN architecture that won the SE track of L3DAS23 challenge. In this paper, we extend the multi-layer distillation within correlated sets to intra-inter set global knowledge transfer, while implementing time-frequency cross-assignment weighting in the distillation loss. The backbone architecture remains DPDCRN due to its facility for structural compression and scalability to both single-channel and multi-channel SE tasks. The backbone architecture of the teacher model is presented in Fig. 1. The main structure of DPDCRN consists of a convolutional encoder, a convolutional decoder, and an intermediate frequency-time (F-T) processing module. The encoder contains two convolutional layers and four dilated convolutional blocks, while the decoder contains two deconvolutional layers and four dilated convolutional blocks. Each frequency-time processing module contains dual branches for frequency and temporal axes, with each branch comprising a causal multi-head self-attention (MHSA) module and a feedforward network based on gated recurrent units (GRUs). For each convolutional layer, the student has half the number of channels of the teacher. For the F-T processing modules, the number of hidden units in the student is half that of the teacher. Additionally, the teacher model has four cascaded F-T modules, while the student only uses one. Overall, the student contains merely 17% of the teacher’s parameters.
III-B Intra-inter Set Knowledge Transfer Architecture
@@
According to the encoder-decoder framework of current mainstream SE models, we partition all layers of teacher-student networks into three correlated sets corresponding to the encoder, decoder, and intermediate F-T processing modules, respectively, to achieve hierarchical knowledge transfer. In previous work [16], we utilized residual fusion to implement cross-layer knowledge transfer within each correlated set and confirmed its advantages over conventional layer-wise distillation. However, residual fusion distillation introduces redundant replication operations when the number of modules in teacher and student networks differs [16], and cross-set information interaction remains unexplored. In this paper, we propose a unified cross-layer distillation framework that jointly addresses intra-set and inter-set knowledge transfer. For intra-set operations, single-layer student features are aligned with all corresponding teacher representations through weighted matching within the correlated set. Regarding inter-set transfer, we first generate representative fusion features for each set via residual fusion, then perform multi-layer knowledge transfer on the fused feature set to facilitate cross-hierarchical circulation of intermediate representations. The overall architecture is shown in Fig. 2.
III-C Time-frequency Cross-calibrated Multi-layer Matching Distillation
@@
The candidate correlation set of the teacher and student is denoted as ${\cal P}=\left\{{\left({{l_{s}},{l_{t}}}\right)\left|{\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]}\right.,{l_{t}}\in\left[{1,\ldots,{L_{t}}}\right]}\right\}$, where ${{l_{s}}}$ and ${{l_{t}}}$ represent the corresponding layers of the candidate student and teacher, respectively. In the feature distillation paradigm, the feature maps of candidate teacher-student pairs are first transformed into specific embedding representations, and then, the distance between teacher-student embeddings is calculated as the distillation loss ${{\cal L}_{KD}}$:
${{\cal L}_{KD}}=\sum\limits_{\left({{l_{s}},{l_{t}}}\right)\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{{l_{s}}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{{l_{t}}}}}\right)}\right)},$ (1)
@@
where ${{\bf{F}}_{{l_{s}}}}$ and ${{\bf{F}}_{{l_{t}}}}$ denote the hidden layer feature representations of the student and teacher respectively, $Tran{s^{s}}$ and $Tran{s^{t}}$ are used to transform teacher-student feature pairs into specific embedding representations for alignment. Conventional layer-wise distillation imposes strict constraints on structural and layer alignment between teacher and student models, requiring redundant layer information to be discarded in cases of mismatch. In contrast, multi-layer distillation that matches single-layer student features with multi-layer teacher features, is not constrained by the difference in the number of layers.

### v2

@@ Abstract
In recent years, complexity compression of neural network (NN)-based speech enhancement (SE) models has gradually attracted the attention of researchers, especially in scenarios with limited hardware resources or strict latency requirements. In this paper, we propose an intra-inter representation fusion knowledge distillation (KD) framework with time-frequency calibration (I2RF-TFCKD) for SE, which achieves distillation through the fusion of multi-layer teacher-student feature flows. Different from previous distillation strategies for SE, the proposed framework fully utilizes the time-frequency differential information of speech while promoting global knowledge flow. Firstly, we construct a collaborative distillation paradigm for intra-set and inter-set correlations. Within a correlated set, multi-layer teacher-student features are pairwise matched for calibrated distillation. Subsequently, we generate representative features from each correlated set through residual fusion to form the fused feature set that enables inter-set knowledge interaction. Secondly, we propose a multi-layer interactive distillation based on dual-stream time-frequency cross-calibration, which calculates the teacher-student similarity calibration weights in the time and frequency domains respectively and performs cross-weighting, thus enabling refined allocation of distillation contributions across different layers according to speech characteristics. The proposed distillation strategy is applied to the dual-path dilated convolutional recurrent network (DPDCRN) that ranked first in the SE track of the L3DAS23 challenge. To evaluate the effectiveness of I2RF-TFCKD, we conduct experiments on both single-channel and multi-channel SE datasets. Objective evaluations demonstrate that the proposed KD strategy consistently and effectively improves the performance of the low-complexity student model and outperforms other distillation schemes. The codes are available at https://github.com/JMCheng-SEU/I2RF-TFCKD-SE.
keywords:
[…]
1 Introduction
@@
Speech enhancement (SE) aims to remove background noises and recover clean speech from noisy mixtures. As a perception front-end algorithm, its applications have already penetrated into diverse domains including such as intelligent terminals, industrial manufacturing, medical health, etc. [40]. With the recent emergence of data-driven deep learning approaches, SE models have begun leveraging massive training data to characterize both speech signals, noise components, and their nonlinear relationships. Capitalizing on the representational learning capabilities of deep neural networks (DNNs), deep learning-based SE methods have achieved superior noise suppression performances and have attained state-of-the-art results in international SE challenges [24], gradually becoming the mainstream SE solutions.
@@
Model compression is an effective solution, including pruning, quantization, and knowledge distillation (KD). Tan et al. [27] reduce the complexity of SE networks using three different techniques: sparse regularization, iterative pruning, and clustering-based quantization. However, pruning and quantization still heavily depend on the original model structure and hardware support. In contrast, KD transfers knowledge from a powerful teacher to a compact student, offering better generality and flexibility—thus becoming the technical path chosen in this paper. Actually, many researchers have already focused on distillation methods for SE models. Kim et al. [13] use the teacher’s pseudo target to train the student model, enabling a zero-shot learning procedure. Thakker et al. [28] integrate KD with multi-task learning losses from speech recognition to achieve SE model compression. These studies primarily focus on minimizing output-level discrepancies between teachers and students while neglecting intermediate feature representations. Recently, Wan et al. [31] propose a hierarchical attention-based distillation framework for SE, exploring the utilization of intermediate representations. Similarly, Nathoo et al. [21] introduce a fine-grained similarity-preserving KD loss, which aims to match the student’s intra-activation Gram matrices to that of the teacher. Knowledge transfer of intermediate model representations has gradually attracted researchers’ attention, with the key challenge lying in the SE-specific distillation design.
@@
In this paper, we design an intra-inter representation fusion with time-frequency calibration KD (I2RF-TFCKD) framework to further optimize the cross-layer supervision distillation paradigm. On the one hand, unlike previous approaches [21, 4] that conduct cross-layer interactions within isolated correlation sets, we propose collaborative distillation across intra-set and inter-set correlations, effectively promoting knowledge flow throughout the model architecture. On the other hand, we introduce temporal and spectral domain cross-computation to derive multi-layer distillation calibration weights, leveraging time-frequency characteristics for targeted distillation optimization. The backbone network for distillation employs our previously proposed dual-path dilated convolutional recurrent network (DPDCRN) model [5], which won the SE track of the L3DAS23 challenge. The main contributions of this work are as follows:
@@
  • We design a collaborative distillation paradigm for intra-set and inter-set correlations. Specifically, the SE model is partitioned into multiple correlated sets, and multi-layer teacher–student matching distillation is performed both within each set and across sets.
  • We introduce a recursive fusion mechanism to generate representative features for each correlated set from both teacher and student, and then integrate them into a fused feature set. The recursive fusion preserves the critical information of each set while reducing redundancy in cross-set pairing.
  • We propose multi-layer time-frequency cross-calibration, which independently computes teacher-student similarity calibration weights along temporal and spectral dimensions for cross-weighting, enabling refined allocation of distillation contributions across different layers.
@@
  • Extensive evaluations on public SE benchmarks demonstrate that our proposed I2RF-TFCKD framework for SE surpasses existing distillation methods and enabling the student model to achieve competitive enhancement results with low computational complexity.
The remainder of the paper is organized as follows: Section 2 introduces related works. Section 3 presents a detailed description of the proposed framework. The experimental setup is outlined in Section 4, while Section 5 covers the experimental results and discussion. Finally, conclusions are presented in Section 6.
2 Related work
2.1 Time-frequency processing for SE
Speech signals inherently exhibit temporal evolution characteristics (e.g., syllabic rhythm and prosody) and frequency properties (e.g., formant and harmonic structures). Time-domain processing captures the nonuniform variations of speech energy along the temporal axis, whereas frequency-domain representations are well suited to differentiating the distribution patterns of speech and noise. Consequently, the joint processing of time and frequency domains (TF processing) has received increasing attention in recent SE research [40].
On one hand, TF processing has been deeply integrated into the architectural design of SE networks. Models based on convolutional–recurrent frameworks [12, 39] utilize convolutional modules to extract local spectral patterns in the frequency domain, while recurrent units are used to capture temporal dependencies in the time domain. More recent designs [14, 22] adopt parallel time–frequency dual-path branches, unfolding intermediate representations sequentially along the time and frequency axes and employing recurrent units to alternately process and integrate information across both domains. All these architectures explicitly emphasize the interaction between time and frequency domains, jointly preserving temporal continuity and spectral details.
On the other hand, TF interaction has also been reflected in the design of loss functions for SE models. Xia et al. [36] introduce two mean-squared-error (MSE)-based spectral losses, where frame-level voice activity detection (VAD) in the time domain is used to separate the speech distortion and noise suppression terms. The modulation-domain loss [30], formulated via the spectro-temporal modulation index (STMI), evaluates the integrity of processed speech within a perceptually motivated TF modulation space. These TF domain losses fully leverage speech-specific characteristics to formulate regression distance measures. Overall, TF processing enables SE models to simultaneously capture temporal context and perceive fine spectral structures, thereby providing a promising direction for optimizing SE frameworks.
2.2 Knowledge distillation for SE
@@
Knowledge distillation [11] has been widely applied in the fields of computer vision and natural language processing. In the acoustic field, KD methods are initially adopted for classification tasks such as automatic speech recognition [37]. Subsequently, KD for regression tasks like text-to-speech (TTS) and SE has gradually emerged [19]. Early distillation studies for SE models primarily focus on minimizing output-level discrepancies between teachers and students, exemplified by elite sub-band distillation [10]. These works aim to align student models with teacher patterns at the learning objective level but overlook the utilization of intermediate representations containing richer information. Feature distillation holds greater potential for SE models as their architectures and functionalities grow increasingly complex. A common challenge in representation-based distillation arises from dimension mismatch between teacher and student features.
To address this, a cross-layer similarity distillation framework [3] is proposed for SE models, which aligns teacher-student features via similarity loss. Similarly, multiple studies have adopted multi-layer distillation strategies to compress SE models, including bidirectional KD [1] and two-step fine-grained similarity-preserving KD [21]. It is worth noting that recent studies have begun to focus on leveraging the inherent characteristics of the SE task to improve distillation frameworks. Dynamic frequency adaptive distillation [38] dynamically evaluates model outputs and adjusts distillation targets according to the requirements of different speech frequency bands. An independent attention transfer mechanism is designed for temporal and channel dimensions in [8] to facilitate rich knowledge transfer across different structured models. However, existing studies have not fully exploited the time-frequency characteristics of speech in designing feature distillation frameworks. Additionally, cross-layer distillation frameworks such as [21, 4] do not consider the global propagation of knowledge. To address these limitations, this work mainly includes two innovative improvements: the time-frequency calibration strategy for distillation and the intra-inter set global knowledge transfer mechanism.
3 Methodology
3.1 Teacher and student network architecture
@@
In this paper, we extend the multi-layer distillation within correlated sets [21, 4] to intra-inter set global knowledge transfer, while implementing time-frequency cross-assignment weighting in the distillation loss. The backbone architecture chooses DPDCRN [5] that won the SE track of the L3DAS23 challenge, due to its facility for structural compression and scalability to both single-channel and multi-channel SE tasks. The backbone architecture of the teacher model is presented in Fig. 1, where $B$ is the batch size, $C$ is the number of output channels, $T$ is the number of speech frames, $D$ is the feature dimension size. The input of the DPDCRN model is the complex spectrum obtained by applying the short-time Fourier transform (STFT) to the noisy speech. The output is the estimated complex ratio mask, which is multiplied with the noisy spectrum in the complex domain, followed by an inverse STFT (iSTFT) to reconstruct the enhanced speech. The main structure of DPDCRN consists of a convolutional encoder, a convolutional decoder, and an intermediate frequency-time (F-T) processing module. The encoder contains two convolutional layers and four dilated convolutional blocks, while the decoder contains two deconvolutional layers and four dilated convolutional blocks. Each frequency-time processing module contains dual branches for frequency and temporal axes, with each branch comprising a causal multi-head self-attention (MHSA) module and a feedforward network based on gated recurrent units (GRUs). Layer normalization (LNorm) and ReLU activation are incorporated between layers to regulate the distribution of feature flows. For each convolutional layer, the student has half the number of channels of the teacher. For the F-T processing modules, the number of hidden units in the student is half that of the teacher. Additionally, the teacher model has four cascaded F-T modules, while the student only uses one. Overall, the student contains merely 17% of the teacher’s parameters.
Figure 2: Overall architecture of the I2RF-TFCKD framework. We present the detailed process of intra-inter set distillation. Fig. 2(a) shows the time-frequency cross-calibration knowledge transfer and the recursive feature fusion across different layers within a single correlated set. Fig. 2(b) demonstrates inter-set distillation achieved via the fused feature set among various correlated sets. Fig. 2(c) visualizes the process of generating fused features by utilizing the features of the current layer and the inherited recursive features.
3.2 Intra-inter representation fusion knowledge transfer architecture
@@
According to the encoder-decoder framework of current mainstream SE models, we partition all layers of teacher-student networks into three correlated sets corresponding to the encoder, decoder, and intermediate F-T processing modules, respectively, to achieve hierarchical knowledge transfer. In previous work [4], residual fusion is utilized to implement cross-layer knowledge transfer within each correlated set and confirmed its advantages over conventional layer-wise distillation. However, residual fusion distillation introduces redundant replication operations when the number of modules in teacher and student networks differs, and cross-set information interaction remains unexplored. In this paper, we propose a unified cross-layer distillation framework that jointly addresses intra-set and inter-set knowledge transfer. For intra-set operations, single-layer student features are aligned with all corresponding teacher representations through weighted matching within the correlated set. Regarding inter-set transfer, we first generate representative fusion features for each set via residual fusion, then perform multi-layer knowledge transfer on the fused feature set to facilitate cross-hierarchical circulation of intermediate representations. The overall architecture is shown in Fig. 2.
3.2.1 Intra-set multi-layer teacher–student feature distillation
@@
The candidate teacher-student correlation set is denoted as ${\cal P}=\left\{{\left({{l_{s}},{l_{t}}}\right)\left|{\forall{l_{s}}\in\left[{1,\ldots,{L_{s}}}\right]}\right.,{l_{t}}\in\left[{1,\ldots,{L_{t}}}\right]}\right\}$, where $l_{s}$}} and $l_{t}$}} represent the corresponding layers of the candidate student and teacher, respectively. In the feature distillation paradigm, the feature maps of candidate teacher-student pairs are first transformed into specific embedding representations, and then, the distance between teacher-student embeddings is calculated as the distillation loss. As shown in Fig. 3(a), single-layer feature distillation only transfers knowledge between teacher and student at the same hierarchical level:
${{\cal L}_{singleKD}}=\sum\limits_{i\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{l_{s}^{i}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{l_{t}^{i}}}}\right)}\right)},$ (1)
@@
where ${\bf{F}}_{{l_{s}}}$} and ${\bf{F}}_{{l_{t}}}$} denote the hidden layer feature representations of the student and teacher respectively, $Tran{s^{s}}$ and $Tran{s^{t}}$ are used to transform teacher-student feature pairs into specific embedding representations for alignment. $\cal D$} is the computation of distillation distance. However, layer-wise distillation imposes strict constraints on structural and layer alignment between teacher and student models, requiring redundant layer information to be discarded in cases of mismatch. In addition, one-to-one knowledge transfer makes it difficult for the student model to progressively learn hierarchical information from the teacher. Therefore, we introduce multi-layer matching distillation for knowledge transfer within correlated sets. As illustrated in Fig. 3(b), The single-layer student feature is aligned with hierarchical teacher features within the correlated set:
${{\cal L}_{multiKD}}=\sum\limits_{\left({{l_{s}},{l_{t}}}\right)\in{\cal P}}{{\cal D}\left({Tran{s^{s}}\left({{{\bf{F}}_{{l_{s}}}}}\right),Tran{s^{t}}\left({{{\bf{F}}_{{l_{t}}}}}\right)}\right)}.$ (2)
Nevertheless, cross-layer distillation may incur redundant computations. To address this, we assign distillation weights across layers according to semantic information, which will be further discussed in Section 3.3.
Figure 3: (a) using single-level teacher knowledge to guide one-level learning of the student. (b) using multiple layers of the teacher to supervise one layer in the student.
3.2.2 Inter-set distillation based on recursive representation fusion
${{{\bf{\tilde{R}}}}^{j-1}}={\rm{2DConv}}\left({{{\bf{R}}^{j-1}}}\right),$ (3)
${{{\bf{\tilde{F}}}}^{j}}={\rm{UpSampling}}\left({{{\bf{F}}^{j}}}\right),$
${{\bf{T}}^{j}}={\rm{2DConv}}\left({{{\bf{R}}^{j}}}\right).$ (6)
For each correlated set, we select the last recursion output as the representative fused feature of the current set, which is then integrated to form a fused feature set. We believe that due to the symmetry of the encoder-decoder architecture in SE models, higher-level features from the middle layers exhibit stronger capability to learn effective information from lower-level features near the input/output ends. Therefore, in terms of residual fusion direction, the encoder and intermediate F-T processing modules follow a forward order starting from the first layer, while the decoder adopts a reverse order integrating from the final layer towards the middle. Corresponding representative features are generated for each correlated set in both the teacher and student models following the above process, and integrated into the teacher-student fused feature set for inter-set distillation.
3.3 Time-frequency cross calibration for multi-layer hierarchical matching